OpenStudy (anonymous):

use the substitution method to solve 3x+2y=11 and x-2y=9

OpenStudy (anonymous):

i can help you and you can try and help me

OpenStudy (anonymous):

sounds good i don't know where to begin really

OpenStudy (anonymous):

well first covert x-2y=9 to the formula y=mx+b

OpenStudy (anonymous):

convert sorry

OpenStudy (anonymous):

ok so 2y=x=9?

OpenStudy (anonymous):

not plus +

OpenStudy (anonymous):

not exactly well more like subtract x from both sides more like 2y=9-x and divide by two so you can have y alone

OpenStudy (anonymous):

it would be y=-x+9/2 i think

OpenStudy (anonymous):

then plug in y for the y in the paragraph so the equation would be 3x+2(-x+9/2)=11

OpenStudy (anonymous):

then solve

OpenStudy (anonymous):

Ok thank you!